im trying to figure out the equation to determine the shipping cost.
AaBbCcD AaBbCD AaBbCr AaBbCct A 1 Normal 1 No Spac... Heading1 Heading 2 Paragraph Styles Consider the following data that you have obtained regarding the shipping of the item. Number of Blenders 10 20 50 100 200 Shipping Cost (dollars) 38 70 154 254 304 Determine the equation of a quadratic function that best fits the above data. This will be your shipping a. cost function C(x). Use the given instructions to produce the equation and the graph in Excel. Include these in your submission. b. How much will it cost to ship the number of blenders that yields the company its maximum profit? How much will it cost the company to ship the number of blenders that yields the company a profit of 0? Assuming it costs $20 to produce each blender, complete the following chart and calculate the revenue. Fill in the empty rows with data obtained for the number of blenders that yield maximum profit and a profit of zero. c. Number of Production Cost Shipping Cost Profit (dollars)Revenue Blenders(dollars) 10 30 60 100 dollar1314.30 38 (dollars) 200 Do the revenue amounts make sense? If not, what could possibly be a reason? What other factors could be influencing the profit? In 3-5 sentences elaborate on your findings. d. Based on your market research, how many blenders would you propose to the firm that they make? Give your reasoning in at least 1 well developed paragraph. in your suggestions make sure you discuss at least two options that the firm can consider 4. . Include information about the type of area in which you are looking to market this product: a small town, a large mall, online retailer? Make sure to keep in mind that you have to pay for the production of all these blenders and make a reasonable profit.
project [Read-Onlyl- Word ut References Mailings Review View Desin Layout 1 Normal 1 No Spac.. Heading1 Heading 2 Ti nt Paragraph Styles Market Kesearch Project Suppose you are a Market Analyst working for a small appliance manufacturer. Your firm is designi and launching sales of a new blender. Before your firm will actually represent and manufacture the product, you need to determine if it will be in the firm's best financial interest to do this. 1. In the following questions, you will do some research and planning for the launch of your new blender. In order to launch your new appliance business, you need to first convince your investors with some market research. There are few considerations you should make. In four well developed paragraphs using the highlighted vocabulary, respond to the following questions. How do production costs impact your profit? Will advertising also impact your profit margins? Identify and describe other costs associated with trying to produce, market, and sell a product? How will these impact your profit? Will it always be the case that the more of an item you produce and sell, the more profit you will make? How do businesses determine their breakeven point and calculate losses? Shipping of goods often depends upon the weight and the dimensions of the container that is being transported to various stores. How might this impact the number of blenders produced and shipped, as well as the size of the blender? Why might it not be a good idea to overproduce a product? How can this impact the profit? How does sales . price and expected revenue impact the profit? 2. Suppose you find that a firm produces and sells x blenders annually, with their total profit (in hundreds of dollars) given by P (x) 1.11x- 23.943-0.003x2 a) Using a graphing utility produce the graph of P in such a way that the appropriate features of the graph are displayed. Print this graph and include it with your submission. In 1-3 sentences explain how you chose what features of the graph you are attempting to display and why? b) Does the graph of ever dip below the x-axis? In 1-3 sentences explain what this cou ld indicates using the vocabulary from part 1. c) Using the graph obtained in part a) estimate the following in interval notation i. For what values of x will P be negative? l. For what values of x will P be positive?
Answer #1

Standand equahon tay a uadrato equahon is ax2+bx y n 2 VD the 一22. given data Points are (10,38) (2 o, 7ο) .(50,154) (too,254) poo, 304)a016 DC2 1022 5100 +200 10600 5 = 182 6800 = 341284000 5 33+70+154+454+304- y = 6 4 IG 5 19136 2. 2 2. 30 8336 0Sxx = 10600-764 = 4 82 4 Sxy = 19136-ㄱ6x164 6672 SX2y2ニ3412 84000-40600): 229924000 5x3o33360 lo6oo 16 1349 0 a= ta84960や4824